OpenStudy (anonymous):

how do i find an equation of the line ? so it needs to be parallel to y=2x+3, through the point 3,1. how do i do this?

OpenStudy (anonymous):

Parallel means that the k values are the same. y=2x+3, here the k value is 2. We know the or line has to look like this: y = 2x + c We also know that it goes through the point 3,1. We replace the x and the y 1 = 2*3 + c 1 = 6 + c -5 = c Now: y = 2x - 5 http://www.wolframalpha.com/input/?i=y+%3D+2x+-+5+at+3

OpenStudy (anonymous):

say that the eqn is y=mx+c Since it is // to y=2x+3 m=2 Applying (3,1) 1=2*3+c c=-5 y=2x-5

OpenStudy (anonymous):

so you found c then placed it in the original equation?

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

thanks. is that just for when they are parallel that you use the plus c method? what is the equation for this method so i can use it on other problems?

OpenStudy (anonymous):

what if it is perpendicular and what is the equation please?

OpenStudy (anonymous):

If y=2x+3 is given as perpendicular to the line then ur m will b -1/2

OpenStudy (anonymous):

If two lines are perpendicular then m1*m2=-1 If two lines are parallel then m1=m2

OpenStudy (anonymous):

thanks